A sinusoidal travelling wave in x-direction has amplitude of 15 π‘π‘š, wavelength of 40 π‘π‘š and the

frequency of 8 𝐻𝑧. The vertical displacement at π‘₯ = 0 and 𝑑 = 0 is 15 π‘π‘š. Find (i) Wave number (ii)

Time period (iii) Wave speed, and (iv) Angular frequency of wave.


1
Expert's answer
2022-03-07T10:19:17-0500

1.

"k=\frac{2\pi}{\lambda}=15.7~m^{-1},"

2.

"T=\frac 1{\nu}=0 .13~s,"

3.

"v=\lambda\nu=3.2~\frac ms,"

4.

"\omega=2\pi \nu=50.3~\frac{rad}s."